OpenStudy (anonymous):

i really got lost with the other answer somone help me plz.. The point (6, 8) lies on the line 8x + 11y = 136. Determine if this point is a solution of the system 8x + 11y = 136 0x – y = -8

OpenStudy (anonymous):

plug it back into both of those and see if it checks out for both

OpenStudy (anonymous):

8(6) + 11(8) = 136 48 + 88 = 136 check

hero (hero):

Why does your questions always have one of the variables as 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0x - y = -8 0(6) - (8) = -8 0-8 = -8 check

OpenStudy (anonymous):

so yes?

hero (hero):

gj